OpenStudy (czesc):

Giving medal immediately after answer. Find the x-coordinates where f '(x) = 0 for f(x) = 2x + sin(2x) in the interval [0, 2π].

OpenStudy (anonymous):

f'(x)=2+2cos 2x=0 1+cos 2x=0 cos 2x=-1=cos(2n+1) pi 2x=(2n+1)pi \[x=\left( 2n+1 \right)\frac{ \pi }{ 2 }\] where n is an integer. n=0,1 x=?